8. The demand equation for a product is given by: P 30-0.102 At what price is demand unitary elastic? (5 marks)
Answer #1

The price elasticity of demand is the change in quantity demanded corresponding to a change in price.

E(p) = (dQ/dP)/(P/Q)

In the given problem, 0.1Q^2 = 30-P

Or, Q^2= 300- 10P

Q= √(300-10P)

Absolute value of dQ/dP= |1/(√(300-10P)* -10 |.. Equation i

We need to know at what price is elasticity unitary.

Or absolute value of E(p) is 1.

P/Q= P/√(300-10P) ... Equation ii

Multiplying equations i and ii and equating to 1

We get 10.P/ (300-10P)= 1

Or p= 15

Please note, we take absolute value of dQ/dP,

So we ignore the negative.